The perimeter is 46 meters and it's length is 2 meters more than it's width. What is its length?
Answer by TeishaRenee(14)

i like these kind of problems. :)
so your equation for perimeter is 2L+2W=P
and you know that the length (L) is 2 more than the width (W) so that equation would look like L=W+2
and you know that the perimeter is 46
so then you plug that into the equation
2(W+2)+2W=46
the W+2 goes where L would be because that is what l equals.
so then you would multiply the 2 by the w and the 2 by the 2 to get rid of the parenthesis.
2*W=2w and 2*2=4
so then change the equation
2w+4+2w=46
subtract the 4 from both sides.
2w+2w=42
combine the W's
4w=42
divide both sides by 4
w=10.5
then you would go back to the original equation and plug in the w=10.5
2L+2W=46
2L+2(10.5)=46
2L+21=46
subtract the 21 from both sides
2L=25
divide both sides by 2
L=12.5
and then plug it into the equation to make sure it works.
2(12.5)+2(10.5)=46
25+21=46***
Length is 12.5 meters and Width is 10.5 Meters.
email me if you have any other questions.
trdl@msn.com